OpenStudy (anonymous):

Identify the center and radius of the circle with the equation (x - 3)2 + (y - 5)2 = 25.

OpenStudy (anonymous):

C = (5, 3) R = 3 C = (3, 5) R = -5 C = (-3, -5) R = 5 C = (3, 5) R = 5

OpenStudy (anonymous):

@GodlyMaths

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (xapproachesinfinity):

Hey @EuniceOB97 if we have \(\large (x-a)^2+(y-b)^2=r^2\) which is the equation of the circle whose center is (a,b) and radius is r

OpenStudy (xapproachesinfinity):

try to see what is a and b are in your question and r as well

OpenStudy (anonymous):

a is 3 b is 5and r is25

OpenStudy (xapproachesinfinity):

not quite you got a and b but r is not quite correct

OpenStudy (anonymous):

5?

OpenStudy (xapproachesinfinity):

correct since 25=5^2

OpenStudy (anonymous):

Okay how do i know which is center ?

OpenStudy (xapproachesinfinity):

i said the center is (a,b) in my equation

OpenStudy (anonymous):

So the answer choice is C?

OpenStudy (anonymous):

a and b is negative ?

OpenStudy (xapproachesinfinity):

well what do you think

OpenStudy (xapproachesinfinity):

nothing is negative pay attention to the equation that i wrote

OpenStudy (anonymous):

C= (3,5) R= 5?

OpenStudy (xapproachesinfinity):

\(\large (x-a)^2+(y-b)^2=r^2\) see here (a,b) is the center r is the radius!

OpenStudy (xapproachesinfinity):

yes that's the correct answer but do you get what i'm saying

OpenStudy (anonymous):

Yes i do . :)

OpenStudy (xapproachesinfinity):

good

OpenStudy (anonymous):

Six posts are evenly spaced along a circular wall encasing a fountain. Anna stands at point A, where she cannot see beyond the only two posts visible to her. What is m<A?

OpenStudy (anonymous):

100o 120o 125o 150o

OpenStudy (xapproachesinfinity):

post your question separately! i don't have enough time actually